OpenStudy (anonymous):

Suppose the amount of money in a bank account with an initial balance of $25,000 earns 1.5 percent interest per year. If the exponential function, y=A*B^X , is used to model this situation, which value would represent B in the function?

OpenStudy (anonymous):

Sorry I have so many questions. I'm taking an exam and I really want to get a good grade. I suck at math.

OpenStudy (anonymous):

1.015 = B 1 for the original 0.15 for the interest

OpenStudy (kash_thesmartguy):

No direct answers!

OpenStudy (anonymous):

Man you are a life saver. Do you think you could help me with another problem? And not give me the answer.

OpenStudy (kash_thesmartguy):

Read the "Terms and Conditions" and "Code of Conduct"

OpenStudy (anonymous):

ok

OpenStudy (kash_thesmartguy):

^^ That was for @peachpi

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

Well can someone help please?

OpenStudy (anonymous):

@Kash_TheSmartGuy it's kind of hard not to give a direct answer to that question. There's nothing to work out

OpenStudy (kash_thesmartguy):

Well, you can give them the formula or explanation of how you got the answer! Don't just give away answers!

OpenStudy (anonymous):

what are you talking about? The question is asking him to pick out a number for use in a formula. And I gave a perfectly good explanation of how to do it

OpenStudy (anonymous):

So can someone help please?

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

well its 2 actually

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so first question: Which exponential function is best represented by this table? x f(x) 0 2 1 6 2 18 3 54 answers: f(x)=3(2)^x f(x)=2(3)^x f(x)=3(1/2)^x f(x)=2(1/3)^x

OpenStudy (anonymous):